OpenStudy (anonymous):

Need help finding the complex solutions for the differential equation y''' + 3y'' +28y' +26y = 0

OpenStudy (anonymous):

Sorry, general solution

OpenStudy (lgbasallote):

if i recall right \[y = y_c\] since it's homogeneous

OpenStudy (anonymous):

well.....I basically need help factoring to get one regular solution and 2 complex solutions.

OpenStudy (anonymous):

and not a clue, sorry

OpenStudy (lgbasallote):

jumping to the good stuff you have \[\huge m^3 +3m^2 + 28m + 26 = 0\] right?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

i know the real root will be negative...

OpenStudy (anonymous):

-1

OpenStudy (lgbasallote):

really?

OpenStudy (anonymous):

so that means i factor out a (m+1)?

OpenStudy (lgbasallote):

hmm yep that's right

OpenStudy (anonymous):

the actual answer is ce^(-x) + ce^(-x)cos5x + ce^(-x)sin5x

OpenStudy (lgbasallote):

so what exactly is your problem?

OpenStudy (anonymous):

so i think that means there are to factorials (wording?) with complex numbers 'i'?

OpenStudy (anonymous):

two*

OpenStudy (anonymous):

i'm looking at alphawolf and it says that they are: -1-5i and -1+5i. do you know how that is so?

OpenStudy (lgbasallote):

divide m^3 + 3m^2 + 28m + 26 by m+1

OpenStudy (lgbasallote):

m^2 + 2m + 26 ______________________________ m+1 | m^3 + 3m^2 + 28m + 26 m^3 + m^2 ------------------- 2m^2 + 28m 2m^2 + 2m ---------- 26m + 26

OpenStudy (lgbasallote):

so the quotient is m^2 + 2m + 26

OpenStudy (lgbasallote):

it's not factorable so you use quadratic formula

OpenStudy (lgbasallote):

\[\huge m = \frac{-2 \pm \sqrt{2^2 - 4(26)}}{2}\] \[\huge \implies m=\frac{-2 \pm \sqrt{4 - 104}}{2}\] \[\huge \implies m=\frac{-2 \pm \sqrt{-100}}{2}\] \[\huge \implies m= \frac{-2 \pm 10i}{2}\] \[\huge \implies m=-2 \pm 5i\]

OpenStudy (lgbasallote):

uhhh that should be \[\huge \implies m = -1 \pm 5i\]

OpenStudy (lgbasallote):

does that help?

OpenStudy (anonymous):

absolutely! this is where my teacher skipped

OpenStudy (lgbasallote):

haha now you know ^_^

OpenStudy (anonymous):

thank you

OpenStudy (lgbasallote):

welcome :D